OpenStudy (anonymous):

Please I'll give you a medal :) what is 6 radical 4?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

or -12

OpenStudy (anonymous):

12

OpenStudy (anonymous):

radical 4 is either 2 or negative two. multiply that by the coefficient (6) and you get 12 or -12

OpenStudy (anonymous):

lol i messed up

OpenStudy (anonymous):

^^and by "either" I mean "both". Radicals are weird like that.

ganeshie8 (ganeshie8):

radical 4 is always positive 2

ganeshie8 (ganeshie8):

so oly 12 works here

OpenStudy (anonymous):

... I don't think so @ganeshie8

OpenStudy (anonymous):

(-2)^2=4, so therefore the square root of 4 is also -2

OpenStudy (anonymous):

Solve for x. 3/x = x/32 96 48 4 radical 6 6 radical 4

ganeshie8 (ganeshie8):

thats not correct @gandalfwiz google a bit u wil find out

OpenStudy (anonymous):

Yes, you could google it, but it wouldn't give you the complete answer.

OpenStudy (anonymous):

thanks what would this answer be Solve for x. 3/x = x/32 96 48 4 radical 6 6 radical 4

OpenStudy (anonymous):

@ganeshie8 Maybe you're thinking of how if you have a negative number under a square root you get an imaginary number? But this is just that the answer is negative, not the term under the square root.

OpenStudy (anonymous):

@emilysheeran17 The answer I got for that was 4 root 6

ganeshie8 (ganeshie8):

I am saying when you take a square root of +ve number, it never gives u a negative number. It gives oly +ve number.

OpenStudy (anonymous):

@gandalfwiz I love you!

ganeshie8 (ganeshie8):

http://www.wolframalpha.com/input/?i=sqrt%284%29

OpenStudy (anonymous):

@ganeshie8 , that 's prety different than what I've learned in math class. You do agree that (-2)^2 equals 4, right?

OpenStudy (anonymous):

THanks @emilysheeran17 :)

ganeshie8 (ganeshie8):

yes, so ?

OpenStudy (anonymous):

@ganeshie8 look at that site and scroll down.

ganeshie8 (ganeshie8):

you're not getting wat im saying at all

ganeshie8 (ganeshie8):

\(x^2 = 4\) => \(x = \pm \sqrt{4}\)

ganeshie8 (ganeshie8):

but, \(\sqrt{4} = 2\) oly. a square root cann never take a negative value.

OpenStudy (anonymous):

I'm not understanding what you're saying. root 4 -2 only? What do you mean by that?

ganeshie8 (ganeshie8):

its a equal sign, do u read it as - sign ha ?

OpenStudy (anonymous):

Yeah, I saw it as a minus sign. But I don't see it as proof that the square root of 4 is only two.

ganeshie8 (ganeshie8):

I didnt intend it to be a proof lol

ganeshie8 (ganeshie8):

I am trying to convey you wat I see. thats all

OpenStudy (anonymous):

Hey guy one more? Six is the geometric mean of 18 and what number?

OpenStudy (anonymous):

I think -6

ganeshie8 (ganeshie8):

6 radical 4 has oly one solution : 12. it cannot have two solutions.

OpenStudy (anonymous):

no thats not a choice :( the options are 2 3 4 or 9

ganeshie8 (ganeshie8):

let x be the other number :- \(6 = \sqrt{18x}\) \(6^2 = 18x\) \(36 = 18x\) solve x

OpenStudy (anonymous):

@gandalfwiz please help me and @ganeshie8 just give up dude

OpenStudy (anonymous):

@emilysheeran17 sorry, I'm not sure then. Maybe I'm finding the algebraic mean lol.

ganeshie8 (ganeshie8):

fine

OpenStudy (anonymous):

@gandalfwiz What is the geometric mean of 3 and 2?

OpenStudy (anonymous):

Well, the way I learned to do means was adding up the numbers and dividing by the quantity of numbers, but if that didn't work for your previous problem I don't know if it will work for this one either.

OpenStudy (anonymous):

My way would give you a mean of 2.5. Is that an answer choice?

OpenStudy (anonymous):

no sorry I wrote the wrong thing its What is the geometric mean of 3 and 21

OpenStudy (anonymous):

I got 12

OpenStudy (anonymous):

Ohhhhh wait I just looked it up.

OpenStudy (anonymous):

You multiply the numbers together and then take the square root if there are 2 terms, the cube root if there are 3 terms, etc.

OpenStudy (anonymous):

That would mean that the geometric sum of 3 and 21 is squareroot 63, which can be simplified to 3root7